OpenStudy (anonymous):

Explain in your own words how you can simplify x^4/x^2 without having to write the expanded form.

OpenStudy (pradipgr817):

x^(4-2)=x^2

Parth (parthkohli):

There's one formula, which is the following:\[\large{a^m \over a^{n}} = {a^{m - n}} \]

Parth (parthkohli):

Therefore,\[\large{x^4 \over x^2} = {x^{4 - 2}} \]

OpenStudy (anonymous):

can you tell me what that method is called?